1. Express each number as a product of its prime factors:

(i) 140

(ii) 156

(iii) 3825

(iv) 5005

(v) 7429

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solutions: (i) 140 By Taking the LCM of 140, we will get the product of its prime factor. Therefore, 140 = 2 × 2 × 5 × 7 × 1 = 22×5×7 (ii) 156 By Taking the LCM of 156, we will get the product of its prime factor. Hence, 156 = 2 × 2 × 13 × 3 × 1 = 22× 13 × 3 (iii) 3825 By Taking the LCM of 3825, we will get the product of its prime factor. Hence, 3825 = 3 × 3 × 5 × 5 × 17 × 1 = 32×52×17 (iv) 5005 By Taking the LCM of 5005, we will get the product of its prime factor. Hence, 5005 = 5 × 7 × 11 × 13 × 1 = 5 × 7 × 11 × 13 (v) 7429 By Taking the LCM of 7429, we will get the product of its prime factor. Hence, 7429 = 17 × 19 × 23 × 1 = 17 × 19 × 23